The chart to the right shows a country's annual egg production. Model the data in the chart with a linear function, using the points (1994,52.7) and (1998,61.2). Let x
represent the year, where x = 0 represents 1994, x= 1 represents 1995, and so on, and let y represent the egg production (in billions). Predict egg production in 2000.
The linear model for the data is
(Type your answer in slope-intercept form. Use integers or decimals for any numbers in the equation. Round to three decimal places as needed.)
Year
1994
1995
1996
1997
1998
1999
2000
Egg products (in billions)
52.7
53.6
55.4
58.1
61.2
64.8
70.7
Check answer

Answer by math_tutor2020(3817)

Answers:
Regression line equation is roughly y = 2.936x + 50.693
The predicted production level in the year 2000 is approximately 68.3 billion

Explanation:

I used GeoGebra's FitLine command to find the regression line.
You could compute this by hand, but I recommend using software to make your life easier.
In real world settings, you'll be using software to speed up the process and ensure no errors happen.

Once we have the regression line equation, plug in x = 6 to represent the year 2000 (since 2000-1994 = 6)
So,
y = 2.936x + 50.693
y = 2.936*6 + 50.693
y = 68.309
y = 68.3
We predict about 68.3 billion eggs are produced in the year 2000
The actual production level is 70.7 billion according to the table.
This is an error of about 70.7 - 68.3 = 2.4 billion

Side note: the correlation coefficient is roughly r = 0.9704